# Transparent and Coherent Procedural Mistake Detection

### v2

@@ D.4 Stopping Criteria Details
The stopping criteria hyperparameter $n$ is fixed at a value of 10. We chose to fix $n$ because tuning $n$ prevents objective comparison of the number of iterations taken by each approach, and $\delta$ and $\epsilon$ can still control the number of iterations of VLMs’ self-dialog. The hyperparameters $\delta$ and $\epsilon$ are selected from a grid search on the validation data over combinations of $\delta\in\{0.05,0.1,0.2,0.4\}$ and $\epsilon\in\{0.025,0.05,0.1,0.2\}$. We maximize overall accuracy, relevance of questions, and potential informativeness of their answers. Specifically, we use a cascading summary metric which is the product of example-level informativeness (reference-adjusted) and example-level relevance if the VLM makes a correct mistake detection classification and this product is positive, else it is zero. Accuracy depends on the mistake confidence threshold $\tau$, which is selected as described in the previous subsection. Selected stopping criteria hyperparameter values for the results in Table 2 are listed later in Appendix E.6, while values for the results in Table 2 are listed in Appendix E.3.
@@ E.2 DPO Training Data Composition Ablation
E.3 Question Generation Fine-Tuning and Inference Details
[…]
@@ E.3 Question Generation Fine-Tuning and Inference Details
Table 6: Selected training data generation hyperparameters for the results presented in Tables 2, 5, and 17.
@@
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
@@
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
@@
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
@@
Table 7: Selected training and inference hyperparameters for the results presented in Tables 2, 5, and 17.
@@
E.4 GPT-4o Evaluation
To better contextualize our results with state-of-the-art proprietary large LMs, we additionally evaluated GPT-4o OpenAI et al. (2024b) [^23] Specifically, we use the August 6, 2024 version of GPT-4o available in Azure AI Foundry. on our Ego4D-PMD dataset for coherent PMD. While proprietary models like GPT-4o offer limited customization, making many of the experiments we presented for open-source VLMs, evaluating vanilla GPT-4o serves as a reference point. It is worth noting that even accessed through APIs, GPT-4o returns responses too slowly to be viable for online, frame-by-frame use.
@@ E.4 GPT-4o Evaluation
The inference hyperparameters $n$, $\delta$, $\epsilon$, and $\tau$ are selected as in our open-source model results, and listed in Table 9. The results on the validation data are listed in Table 10, while the results on the testing data are listed in the main body of the paper in Table 2. Comparing GPT-4o to the base VLMs we evaluated, it is generally inferior in PMD accuracy and informativeness, but asks more relevant questions, runs for fewer iterations, and has higher information gain. Our best model configurations, though, outperform GPT-4o under all evaluation metrics except information gain. [^24] LLaVA with coherence-based fine-tuning, coherence-based re-ranking, and in-context learning achieves .742 bits of information gain, while GPT-4o achieves up to .793 bits. This shows that GPT-4o generally exhibits higher confidence despite having much lower PMD accuracy than our approaches, which is not necessarily an advantage. This suggests that GPT-4o is a reasonable starting point for coherent PMD, but like other off-the-shelf VLMs we evaluated, it may require additional interventions (e.g., coherence-based ranking) to be viable for this task.
$n$ $\delta$ $\epsilon$ $\tau$
@@
E.5 Rationale-Free Evaluation
@@
Interestingly, however, the rationale-free approach achieves generally better accuracy than the coherent PMD results in Tables 2 and 14 with likelihood-based ranking and no in-context learning. However, upon introducing coherence-based ranking and in-context learning, InstructBLIP and LLaVA achieve better accuracy than in the rationale-free approach. Further, the information gain in the rationale-free setting is consistently lower than those achieved in coherent PMD. This demonstrates that while the added transparency of incorporating rationales into PMD does cost the VLM some accuracy, improving the coherence of these rationales (e.g., through the approaches presented in this work) can recover this accuracy and more while enabling more confident decisions from VLMs. Rationale-free GPT-4o achieves the highest observed accuracy of 69.2%. Nonetheless, this accuracy is still low enough for errors to be common, thus necessitating the generation of a rationale for the user.
@@
Table 11: Selected values of mistake confidence threshold $\tau$ for rationale-free PMD with various VLMs. $n$, $\delta$, and $\epsilon$ are no longer used in rationale-free PMD, but $\tau$ is still tuned as in previous experiments.
@@
E.6 Question Selection Inference Hyperparameters and Validation Results
In Table 13, we list the inference hyperparameters for the question selection results presented in Table 2: number of iterations $n$, early stopping $\delta$ and $\epsilon$, and mistake confidence threshold $\tau$. In Figure 5, we use detection error tradeoff (DET) curves to visualize the range of accuracy achieved with all candidate mistake confidence thresholds $\tau$ for the approaches compared in Table 2. The full validation set results with selected hyperparameters are listed in Table 14.
@@ E.6 Question Selection Inference Hyperparameters and Validation Results
Rank ICL $n$ $\delta$ $\epsilon$ $\tau$
@@
Rank ICL $n$ $\delta$ $\epsilon$ $\tau$
@@
Rank ICL $n$ $\delta$ $\epsilon$ $\tau$
@@
E.7 Analysis of Question Sources in In-Context Learning
@@
E.8 Diversity-Based Ranking Baseline
@@
E.9 DPO with Length Penalty Experiment
@@
While the primary purpose of this work was to recast the problem of PMD into a more transparent formulation and lay a foundation for research toward coherent question generation and answering in PMD, we performed an initial experiment to guide future work along this line. Specifically, we imposed an exponential length penalty $l=-1.0$ to text generation during training data generation and inference for evaluation. [^26] Like the results presented in Table 2, candidates generated through in-context learning are also included in training data. During beam search, a length penalty modifies the total log-likelihood $p_{Q}$ of a partial candidate question $Q$ as follows:
@@
Table 17: Ego4D-PMD test set results for DPO-trained VLMs with an additional length penalty $l=-1$ applied during training data generation and inference. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.10 Using Coherence Metrics to Diagnose Common VLM Behaviors
To deepen the insights from the graphs in Figure 4, in Figure 7, we provide several example outputs from LLaVA with coherence-based ranking, which displays a range of behaviors. Below, we further explain these behaviors and examples.
@@ E.10 Using Coherence Metrics to Diagnose Common VLM Behaviors
Example A: Pick up a sink brush from the kitchen slab. Refer to caption Label: ✓        Predicted: ✓ Rationale: 1. Is the sink brush in the person’s hand? Yes
Example B: Tighten the screw. Refer to caption Label: ✓        Predicted: ✗ Rationale: 1. Is the person wearing gloves? No 2. Is the person wearing protective gear? No 3. Is the person wearing a mask? No
Example C: Paint the stone. Refer to caption Label: ✗        Predicted: ✗ Rationale: 1. Is the person wearing a shirt? Unsure
Example D: Drop the bottle of mustard on the countertop. Refer to caption Label: ✓        Predicted: ✗ Rationale: 1. Is the bottle of mustard on the countertop? Yes 2. Is the bottle of mustard on the floor? Yes
Example E: Fold the cut piece of cloth. Refer to caption Label: ✗        Predicted: ✗ Rationale: 1. Is the person working on a piece of cloth? No
Example F: Put the bolt remover in the lawn tractor. Refer to caption Label: ✗        Predicted: ✗ Rationale: 1. Is the bolt remover in the lawn tractor? Unsure 2. Is the bolt remover inside the lawn tractor? Unsure 3. Is the bolt remover attached to the lawn tractor? Unsure
Example G: Put the trowel in a bin. Refer to caption Label: ✓        Predicted: ✗ Rationale: 1. Is the trowel in a bin? No
Example H: Put the bottle in the cabinet. Refer to caption Label: ✗        Predicted: ✓ Rationale: 1. Is the bottle in the cabinet? Yes
Figure 7: Sample outputs from LLaVA with coherence-based ranking, representing the range of behaviors visualized in Figure 4. Images cropped for visibility and space.

### v3

@@ D.4 Stopping Criteria Details
The stopping criteria hyperparameter $n^{*}$ is fixed at a value of 10. We chose to fix $n^{*}$ because tuning $n^{*}$ prevents objective comparison of the number of iterations taken by each approach, and $\delta$ and $\epsilon$ can still control the number of iterations of VLMs’ self-dialog. The hyperparameters $\delta$ and $\epsilon$ are selected from a grid search on the validation data over combinations of $\delta\in\{0.05,0.1,0.2,0.4\}$ and $\epsilon\in\{0.025,0.05,0.1,0.2\}$. We maximize overall accuracy, relevance of questions, and potential informativeness of their answers. Specifically, we use a cascading summary metric which is the product of example-level informativeness (reference-adjusted) and example-level relevance if the VLM makes a correct mistake detection classification and this product is positive, else it is zero. Accuracy depends on the mistake confidence threshold $\tau$, which is selected as described in the previous subsection. Selected stopping criteria hyperparameter values for the results in Table 2 are listed later in Appendix E.7, while values for the results in Table 2 are listed in Appendix E.4.
@@ E.2 DPO Training Data Composition Ablation
E.3 VLM Self-Dialog Runtime Analysis
As discussed in Section Limitations, as task guidance and PMD will ultimately operate in real-world applications, their responsiveness becomes crucial. While detailed inquiries and methods for more efficient PMD are out of scope for this work, we analyze the runtime of various configurations of LLaVA in Table 6 as information for future work in this area. As expected, coherence-based ranking and in-context learning significantly increase the runtime of models up to about 1 minute per example (not suitable for practical application). Meanwhile, applying the coherence-trained VQG adapter does not have much runtime overhead compared to off-the-shelf LLaVA; both take approximately 3.3 seconds per iteration, and the adapter slightly reduces the total runtime due to taking fewer iterations (more appropriate for practical application).
Rank ICL DPO Mean # Iterations Std. # Iterations Mean Runtime Std. Runtime Mean Runtime/Iteration
Likelihood 3.31 1.56 11.1 5.79 3.30
Coherence 3.03 1.71 30.5 20.8 9.48
Coherence 3.50 2.66 61.3 53.6 16.3
Coherence 1.81 1.61 28.2 28.5 14.9
Likelihood 2.26 1.37 7.74 4.92 3.38
Table 6: Runtime analysis for various configurations of LLaVA without and with coherence-based ranking, in-context learning, and a DPO-trained adapter for coherent question generation. Runtimes are measured in seconds per example or iteration. Models are evaluated on the entire validation set of Ego4D-PMD). To better represent a real-time application, models are evaluated one example at a time rather than in batch, which causes small discrepancies with the number of iterations reported in other validation results.
E.4 Question Generation Fine-Tuning and Inference Details
When generating the training data from a specific self-dialog iteration, data is omitted if only one candidate question was generated, or if the chosen question has an unsure answer from the VLM (based on the sureness threshold of 60%). To maximize training data quality, the inference hyperparameters for training data generation are selected based on the training data using the procedure described in Appendix E.7. The VQG adapter is trained for 10 epochs with a batch size maximized for our available GPU memory (4). The learning rate $\eta$ and DPO $\beta$ hyperparameters are selected from a grid search over combinations of $\eta\in\{\text{1e-6},\text{2.5e-6},\text{5e-6},\text{7.5e-6},\text{1e-5}\}$ and $\beta\in\{0.05,0.1,0.5\}$, minimizing the minimum validation set DPO loss (as defined by Rafailov et al., 2023) achieved across all epochs. In each run, the learning rate is warmed up to its assigned value for the first 5% of training steps, then linearly decreased to zero through the remaining steps. Selected hyperparameters for training data generation are listed in Table 7, while selected hyperparameters for training and inference are listed in Table 8. Training is distributed across 4 A40 GPUs, and takes up to about 12 hours. The full validation set results for the results in Tables 2, 5, and 18 with corresponding selected hyperparameters are listed in Table 9. Note that testing results for DPO without in-context learning in training data generation are introduced earlier in Appendix E.2, while those for DPO with length penalty in training data generation are introduced later in Appendix E.10.
@@ E.4 Question Generation Fine-Tuning and Inference Details
Table 7: Selected training data generation hyperparameters for the results presented in Tables 2, 5, and 18.
@@
Rank ICL $\eta$ $\beta$ $n^{*}$ $\delta$ $\epsilon$ $\tau$
@@
Rank ICL $\eta$ $\beta$ $n^{*}$ $\delta$ $\epsilon$ $\tau$
@@
Rank ICL $\eta$ $\beta$ $n^{*}$ $\delta$ $\epsilon$ $\tau$
@@
Table 8: Selected training and inference hyperparameters for the results presented in Tables 2, 5, and 18.
@@
E.5 GPT-4o Evaluation
To better contextualize our results with state-of-the-art proprietary large LMs, we additionally evaluated GPT-4o OpenAI et al. (2024a) [^26] Specifically, we use the August 6, 2024 version of GPT-4o available in Azure AI Foundry. on our Ego4D-PMD dataset for coherent PMD. While proprietary models like GPT-4o offer limited customization, making many of the experiments we presented for open-source VLMs, evaluating vanilla GPT-4o serves as a reference point. It is worth noting that even accessed through APIs, GPT-4o returns responses too slowly to be viable for online, frame-by-frame use.
@@ E.5 GPT-4o Evaluation
The inference hyperparameters $n^{*}$, $\delta$, $\epsilon$, and $\tau$ are selected as in our open-source model results, and listed in Table 10. The results on the validation data are listed in Table 11, while the results on the testing data are listed in the main body of the paper in Table 2. Comparing GPT-4o to the base VLMs we evaluated, it is generally inferior in PMD accuracy and informativeness, but asks more relevant questions, runs for fewer iterations, and has higher information gain. Our best model configurations, though, outperform GPT-4o under all evaluation metrics except information gain. [^27] LLaVA with coherence-based fine-tuning, coherence-based re-ranking, and in-context learning achieves .742 bits of information gain, while GPT-4o achieves up to .793 bits. This shows that GPT-4o generally exhibits higher confidence despite having much lower PMD accuracy than our approaches, which is not necessarily an advantage. This suggests that GPT-4o is a reasonable starting point for coherent PMD, but like other off-the-shelf VLMs we evaluated, it may require additional interventions (e.g., coherence-based ranking) to be viable for this task.
$n^{*}$ $\delta$ $\epsilon$ $\tau$
@@
E.6 Rationale-Free Evaluation
@@
Interestingly, however, the rationale-free approach achieves generally better accuracy than the coherent PMD results in Tables 2 and 15 with likelihood-based ranking and no in-context learning. However, upon introducing coherence-based ranking and in-context learning, InstructBLIP and LLaVA achieve better accuracy than in the rationale-free approach. Further, the information gain in the rationale-free setting is consistently lower than those achieved in coherent PMD. This demonstrates that while the added transparency of incorporating rationales into PMD does cost the VLM some accuracy, improving the coherence of these rationales (e.g., through the approaches presented in this work) can recover this accuracy and more while enabling more confident decisions from VLMs. Rationale-free GPT-4o achieves the highest observed accuracy of 69.2%. Nonetheless, this accuracy is still low enough for errors to be common, thus necessitating the generation of a rationale for the user.
@@
Table 12: Selected values of mistake confidence threshold $\tau$ for rationale-free PMD with various VLMs. $n^{*}$, $\delta$, and $\epsilon$ are no longer used in rationale-free PMD, but $\tau$ is still tuned as in previous experiments.
@@
E.7 Question Selection Inference Hyperparameters and Validation Results
In Table 14, we list the inference hyperparameters for the question selection results presented in Table 2: maximum number of iterations $n^{*}$, early stopping $\delta$ and $\epsilon$, and mistake confidence threshold $\tau$. In Figure 6, we use detection error tradeoff (DET) curves to visualize the range of accuracy achieved with all candidate mistake confidence thresholds $\tau$ for the approaches compared in Table 2. The full validation set results with selected hyperparameters are listed in Table 15.
@@ E.7 Question Selection Inference Hyperparameters and Validation Results
Rank ICL $n^{*}$ $\delta$ $\epsilon$ $\tau$
@@
Rank ICL $n^{*}$ $\delta$ $\epsilon$ $\tau$
@@
Rank ICL $n^{*}$ $\delta$ $\epsilon$ $\tau$
@@
E.8 Analysis of Question Sources in In-Context Learning
@@
E.9 Diversity-Based Ranking Baseline
@@
E.10 DPO with Length Penalty Experiment
@@
While the primary purpose of this work was to recast the problem of PMD into a more transparent formulation and lay a foundation for research toward coherent question generation and answering in PMD, we performed an initial experiment to guide future work along this line. Specifically, we imposed an exponential length penalty l=1.0l=-1.0 to text generation during training data generation and inference for evaluation. [^29] Like the results presented in Table 2, candidates generated through in-context learning are also included in training data. During beam search, a length penalty modifies the total log-likelihood $p_{Q}$ of a partial candidate question $Q$ as follows:
@@
Table 18: Ego4D-PMD test set results for DPO-trained VLMs with an additional length penalty l=1l=-1 applied during training data generation and inference. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.11 Using Coherence Metrics to Diagnose Common VLM Behaviors
To deepen the insights from the graphs in Figure 4, in Figure 5, we provided several example outputs from LLaVA with coherence-based ranking, which displays a range of behaviors. Below, we further explain these behaviors and examples.